# Spec-Gaussian: Anisotropic View-Dependent Appearance for 3D Gaussian Splatting

### v1

@@ Appendix A More Results
A.1 NeRF Synthetic Scenes
@@
As shown in Tabs. 5-7, our method demonstrates the best rendering quality metrics in almost every scene. It’s important to note that the experimental setup for Tri-MipRF [17] differs from other methods. It uses both the training and validation sets as training data, expanding the scale of the model’s data. When its training data is limited to the training set, its metrics suffer a noticeable drop. Nevertheless, to ensure that the experimental results fully reflect the highest performance of each method, and to prevent significant drops in metrics due to differences in experimental environments, we still present the metrics from the Tri-MipRF official paper. Our method achieved more prominent metrics in scenes with notable specular reflection and anisotropy, such as Drums, Lego, and Ship. This demonstrates that our method not only improves the overall rendering quality but also has a more significant advantage in complex specular scenarios.
A.2 NSVF Synthetic Scenes
@@
The NSVF [28] dataset, in comparison to NeRF, features more noticeable metallic specular reflection, as presented in the Wineholder, Steamtrain, and Spaceship scenes. It also includes more complex transmission scenarios, such as Lifestyle. It is important to note that Tri-MipRF fails to converge in the Steam scene with the official code, so we did not report metrics for that scenario. As shown in Tabs. 8-10, we present the per-scene experimental results of PSNR, SSIM, and LPIPS in the supplementary material. The experimental results indicate that compared to other methods based on 3D-GS [21], our method has significant advantages in metallic highlights and complex transmission scenarios. Additionally, we compared it with the SOTA NeRF-based methods based on NeRF. Our approach enables 3D-GS to surpass the latest SOTA of NeRF, achieving high-frequency highlight modeling that 3D-GS couldn’t realize but NeRF could, thereby achieving truly high-quality rendering.
A.3 Anisotorpic Synthetic Scenes
”Anisotropic Synthetic” is a synthetic dataset we rendered ourselves, which includes 8 scenes with significant anisotropy. We tested some existing 3D-GS-based methods on ”Anisotropic Synthetic.” As shown in Tabs. 11-13, our method achieved a very significant improvement in rendering metrics. Qualitative experiments also demonstrate the significant visual advantages of our method, highlighting the substantial improvement our method brings to anisotropic parts, thereby enhancing the overall rendering quality.
A.4 Mip-360 Scenes
The MipNeRF-360 scenes include five outdoor and four indoor scenarios. There are several scenes rich in specular reflections, such as bonsai, room, and kitchen. As shown in Tabs. 14-16, our method achieved significant advantages in the four indoor scenes. This reflects our method’s strengths in modeling specular reflections and anisotropy. In outdoor scenes, our method also achieved rendering metrics comparable to the SOTA methods. Furthermore, with the help of the coarse-to-fine training mechanism, our method significantly reduced the number of floaters, resulting in a substantial improvement in visual effects.
A.5 Tanks & Temples and Deep Blending Scenes
To more comprehensively demonstrate the superiority of our method over 3D-GS in terms of rendering quality, we also compared it with the Deep Blending and Tanks & Temples datasets, which were also used in the original 3D-GS paper. As shown in Tab. 17, our method achieved the best metrics in almost all scenes. This also showcases the versatility of our method, indicating that it is not limited to modeling anisotropic and specular scenarios.
@@ A.5 Tanks & Temples and Deep Blending Scenes
Ours-w anchor 35.57 26.58 35.71 38.12 36.62 30.66 36.81 31.63 33.96
Ours 35.68 26.92 36.14 38.28 36.07 30.85 37.12 31.89 34.12
@@
Ours-w anchor 0.986 0.953 0.987 0.985 0.982 0.962 0.992 0.904 0.969
@@
Ours-w anchor 0.013 0.038 0.012 0.022 0.016 0.037 0.007 0.112 0.032
Ours 0.011 0.032 0.011 0.019 0.014 0.032 0.006 0.104 0.028
@@
Ours-w anchor 40.63 35.56 38.95 38.52 39.47 37.98 36.55 34.04 37.71
Ours 41.52 36.13 39.10 39.60 40.02 38.28 37.43 34.73 38.35
@@
Ours-w anchor 0.994 0.979 0.982 0.994 0.993 0.992 0.984 0.975 0.987
@@
Ours-w anchor 0.005 0.027 0.018 0.007 0.007 0.008 0.021 0.025 0.015
Ours 0.004 0.023 0.015 0.005 0.007 0.007 0.017 0.021 0.013
@@
Ours-w anchor 33.53 31.56 50.35 36.14 32.95 38.48 40.10 30.96 36.76
Ours 34.88 30.83 50.51 37.02 32.90 39.45 41.18 31.46 37.28
Table 11: Per-scene PSNR comparison on our ”Anisotropic Synthetic” dataset.
@@
Ours-w anchor 0.985 0.973 0.999 0.974 0.973 0.988 0.984 0.930 0.976
Ours 0.987 0.967 0.998 0.977 0.967 0.990 0.986 0.943 0.977
Table 12: Per-scene SSIM comparison on our ”Anisotropic Synthetic” dataset.
@@
Ours-w anchor 0.022 0.042 0.004 0.039 0.067 0.017 0.084 0.093 0.046
Ours 0.021 0.051 0.009 0.036 0.087 0.014 0.071 0.087 0.047
Table 13: Per-scene LPIPS (VGG) comparison on our ”Anisotropic Synthetic” dataset.
@@
3D-GS 25.08 21.41 27.26 26.62 22.68 31.54 29.04 31.44 32.16
Scaffold-GS 25.05 21.20 27.33 26.49 23.23 32.13 29.44 31.59 32.49
Ours-w/o anchor 25.11 21.31 27.48 26.59 22.63 31.84 30.05 31.91 33.38
Ours 25.12 21.63 27.50 26.61 23.19 32.14 30.11 32.10 33.68
@@
3D-GS 0.746 0.588 0.855 0.769 0.635 0.924 0.913 0.931 0.944
Scaffold-GS 0.738 0.568 0.846 0.754 0.641 0.927 0.914 0.929 0.946
Ours-w/o anchor 0.739 0.584 0.856 0.759 0.631 0.925 0.919 0.933 0.946
Ours 0.744 0.589 0.850 0.758 0.640 0.929 0.917 0.930 0.950
@@
3D-GS 0.245 0.359 0.123 0.242 0.347 0.199 0.184 0.117 0.182
Scaffold-GS 0.266 0.383 0.143 0.276 0.353 0.200 0.195 0.121 0.186
Ours-w/o anchor 0.247 0.361 0.121 0.246 0.349 0.203 0.180 0.115 0.184
Ours 0.251 0.346 0.137 0.258 0.341 0.192 0.184 0.120 0.174
@@
Method Truck Train Dr Johnson Playroom
[…]
Plenoxels 23.22 0.774 0.335 18.93 0.663 0.422 23.14 0.787 0.521 22.98 0.802 0.499
iNGP 23.38 0.800 0.249 20.46 0.689 0.360 28.26 0.854 0.352 21.67 0.779 0.428
Mip-NeRF360 24.91 0.857 0.159 19.52 0.660 0.354 29.14 0.901 0.237 29.66 0.900 0.252
3D-GS 25.42 0.878 0.147 22.01 0.811 0.209 29.21 0.900 0.247 30.09 0.898 0.247
Scaffold-GS 25.77 0.883 0.147 22.15 0.822 0.206 29.80 0.907 0.250 30.62 0.904 0.258
Ours-w/o anchor 25.50 0.878 0.150 22.38 0.813 0.211 29.24 0.902 0.252 30.17 0.900 0.246
Ours 26.25 0.885 0.144 22.90 0.825 0.204 29.89 0.906 0.251 31.00 0.906 0.253
Table 17: Quantitative comparison on Tanks&Temples and Deep Blending dataset.

### v2

@@ Appendix C More Comparisons
C.1 NeRF Synthetic Scenes
@@
As shown in Tabs. 2-4, our method demonstrates the best rendering quality metrics in almost every scene. It’s important to note that the experimental setup for Tri-MipRF [16] differs from other methods. It uses both the training and validation sets as training data, expanding the scale of the model’s data. When its training data is limited to the training set, its metrics suffer a noticeable drop. Nevertheless, to ensure that the experimental results fully reflect the highest performance of each method, and to prevent significant drops in metrics due to differences in experimental environments, we still present the metrics from the Tri-MipRF official paper. Our method achieved more prominent metrics in scenes with notable specular reflection and anisotropy, such as Drums, Lego, and Ship. This demonstrates that our method not only improves the overall rendering quality but also has a more significant advantage in complex specular scenarios.
C.2 NSVF Synthetic Scenes
@@
The NSVF [27] dataset, in comparison to NeRF, features more noticeable metallic specular reflection, as presented in the Wineholder, Steamtrain, and Spaceship scenes. It is important to note that Tri-MipRF fails to converge in the Steam scene with the official code, so we did not report metrics for that scenario. As shown in Tabs. 5-7, we present the per-scene experimental results of PSNR, SSIM, and LPIPS in the supplementary material. The experimental results indicate that compared to other methods based on 3D-GS [20], our method has significant advantages in metallic highlights and complex transmission scenarios. Additionally, we compared it with the SOTA NeRF-based methods based on NeRF. Our approach enables 3D-GS to surpass the latest SOTA of NeRF, achieving high-frequency highlight modeling that 3D-GS couldn’t realize but NeRF could, thereby achieving truly high-quality rendering as shown in Fig. 19.
[…]
[8.0cm] Dataset NeRF Synthetic Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS $\uparrow$ Num.(k) $\downarrow$ 3D-GS 33.32 0.969 0.031 315 295 w/o ASG 34.03 0.969 0.030 175 271 w/o decoup-MLP 33.95 0.970 0.030 217 244 w/o normal 34.10 0.971 0.029 139 238 Full (Light) 34.08 0.970 0.029 148 186 Full 34.19 0.971 0.028 121 237
Figure 11: Ablation on ASG appearance field.
[…]
[8.0cm] Dataset MipNeRF 360 Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS $\uparrow$ Num.(M) $\downarrow$ 3D-GS 27.47 0.812 0.222 115 3.23 w/o ASG field 27.61 0.830 0.184 113 3.21 w/o c2f 28.01 0.823 0.203 28 3.56 w/ anchor 28.14 0.824 0.196 70 - Full (Light) 28.07 0.834 0.183 44 2.52 Full 28.18 0.835 0.176 33 3.10
Figure 12: Ablation on full method.
Mip-NeRF 360 3D-GS Scaffold-GS Ours-w/o Norm Ours GT
Figure 13: Visualization on Mip-NeRF 360 outdoor scenes. Our method achieves robust floater removal by coarse to fine training.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
3D-GS Scaffold-GS Ours GT
Figure 14: More comparisons with baselines. Our method achieves robust floater removal by coarse-to-fine training.
[…]
[6cm]Refer to caption
Figure 15: Illustration of specular highlights and reflections.
[…]
[7cm] Refer to caption Refer to caption Refer to caption GS-Shader Ours GT
Figure 16: Ccomparison on Ref-NeRF dataset.
Refer to caption Refer to caption Refer to caption Refer to caption
3D-GS Scaffold-GS Ours GT
Figure 17: Visualization on Nex [51] dataset.
@@
Ours-w/ anchor 35.57 26.58 35.71 38.12 36.62 30.66 36.81 31.63 33.96
Ours-light 35.69 26.77 36.03 38.25 36.11 30.84 36.95 31.97 34.08
Ours 35.72 26.92 36.10 38.25 36.46 30.98 37.09 31.97 34.19
@@
Ours-w/ anchor 0.986 0.953 0.987 0.985 0.982 0.962 0.992 0.904 0.969
Ours-light 0.987 0.955 0.988 0.985 0.981 0.963 0.993 0.905 0.970
@@
Ours-w/ anchor 0.013 0.038 0.012 0.022 0.016 0.037 0.007 0.112 0.032
Ours-light 0.012 0.035 0.011 0.019 0.017 0.034 0.006 0.101 0.029
Ours 0.011 0.033 0.011 0.018 0.014 0.031 0.006 0.099 0.028
@@
3D-GS Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Teapot Plane Record Ashtray Dishes Headphone Jupyter Lock
Figure 18: Visualization on our "Anisotropic Synthetic" dataset. We show the comparison between our method and 3D-GS across all eight scenes. Qualitative experimental results demonstrate the significant advantage of our method in modeling anisotropic scenes, thereby enhancing the rendering quality of 3D-GS.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
3D-GS Scaffold-GS GS-shader Ours GT
Figure 19: Visualization on NSVF dataset. Our method significantly improves the ability to model metallic materials compared to other GS-based methods. At the same time, our method also demonstrates the capability to model refractive parts, reflecting the powerful fitting ability of our method.
@@
Ours-w/ anchor 40.63 35.56 38.95 38.52 39.47 37.98 36.55 34.04 37.71
Ours-light 41.48 36.11 39.23 39.54 39.89 38.19 37.22 34.59 38.28
Ours 41.67 36.15 39.33 39.65 40.03 38.26 37.43 34.69 38.40
@@
Ours-w/ anchor 0.994 0.979 0.982 0.994 0.993 0.992 0.984 0.975 0.987
Ours-light 0.995 0.982 0.984 0.993 0.994 0.994 0.984 0.977 0.988
@@
Ours-w/ anchor 0.005 0.027 0.018 0.007 0.007 0.008 0.021 0.025 0.015
Ours-light 0.005 0.024 0.015 0.006 0.007 0.007 0.018 0.022 0.013
Ours 0.004 0.022 0.014 0.005 0.007 0.007 0.017 0.021 0.012
@@
Ours-w/ anchor 33.53 31.56 50.35 36.14 32.95 38.48 40.10 30.96 36.76
Ours-light 34.75 31.01 50.30 37.76 33.03 39.39 41.42 31.68 37.42
Ours 35.24 30.95 50.90 38.03 33.04 40.12 41.47 31.86 37.70
Table 8: Per-scene PSNR comparison on our "Anisotropic Synthetic" dataset.
@@
Ours-w/ anchor 0.985 0.973 0.999 0.974 0.973 0.988 0.984 0.930 0.976
Ours-light 0.987 0.967 0.998 0.984 0.970 0.990 0.987 0.948 0.979
Ours 0.988 0.967 0.999 0.985 0.970 0.990 0.987 0.951 0.980
Table 9: Per-scene SSIM comparison on our "Anisotropic Synthetic" dataset.
@@
Ours-w/ anchor 0.022 0.042 0.004 0.039 0.067 0.017 0.084 0.093 0.046
Ours-light 0.021 0.052 0.007 0.022 0.079 0.014 0.076 0.080 0.044
Ours 0.021 0.051 0.005 0.020 0.077 0.013 0.071 0.075 0.042
Table 10: Per-scene LPIPS (VGG) comparison on our "Anisotropic Synthetic" dataset.
@@
3D-GS 25.63 21.94 27.73 27.02 22.79 31.80 29.12 31.61 32.48
Scaffold-GS 25.61 21.74 27.82 26.79 23.38 32.14 29.62 31.81 32.87
Ours-w anchor 25.44 21.36 27.97 26.91 23.23 32.27 30.12 32.04 33.91
Ours-light 25.87 21.81 28.06 27.23 22.48 32.11 29.74 32.09 33.26
Ours 25.90 21.86 28.07 27.25 22.48 32.11 30.12 32.25 33.54
@@
3D-GS 0.778 0.623 0.874 0.784 0.651 0.928 0.916 0.933 0.948
Scaffold-GS 0.773 0.609 0.867 0.774 0.657 0.931 0.919 0.931 0.950
Ours-w anchor 0.775 0.611 0.869 0.775 0.645 0.932 0.920 0.934 0.953
Ours-light 0.795 0.645 0.879 0.795 0.647 0.934 0.920 0.936 0.952
Ours 0.797 0.648 0.881 0.797 0.647 0.935 0.923 0.937 0.953
@@
3D-GS 0.204 0.328 0.103 0.207 0.318 0.191 0.178 0.113 0.173
Scaffold-GS 0.224 0.339 0.112 0.228 0.315 0.182 0.177 0.114 0.174
Ours-w anchor 0.205 0.292 0.110 0.215 0.293 0.185 0.179 0.115 0.166
Ours-light 0.173 0.279 0.097 0.190 0.275 0.182 0.173 0.111 0.168
Ours 0.166 0.263 0.092 0.184 0.269 0.177 0.166 0.108 0.162
@@
C.3 Anisotorpic Synthetic Scenes
"Anisotropic Synthetic" is a synthetic dataset we rendered ourselves, which includes 8 scenes with significant anisotropy. We tested some existing 3D-GS-based methods on "Anisotropic Synthetic." As shown in Tabs. 8-10, our method achieved a very significant improvement in rendering metrics. Fig. 18 shows the comparison between our method and 3D-GS across all eight scenes. Qualitative experiments also demonstrate the significant visual advantages of our method, highlighting the substantial improvement our method brings to anisotropic parts, thereby enhancing the overall rendering quality.
C.4 Mip-360 Scenes
The MipNeRF-360 scenes include five outdoor and four indoor scenarios. There are several scenes rich in specular reflections, such as bonsai, room, and kitchen. As shown in Tabs. 11-13, our method achieved significant advantages in the four indoor scenes. This reflects our method’s strengths in modeling specular reflections and anisotropy. In outdoor scenes, our method also achieved rendering metrics comparable to the SOTA methods. Furthermore, with the help of the coarse-to-fine training mechanism, our method significantly reduced the number of floaters as shown in Fig. 16, resulting in a substantial improvement in visual effects.